OpenStudy (firejay5):

Graph each function. Identify the domain and range. f(x) = {x/3, x < or equal to 0 2x - 6, 0 < x < 2 1, x > or equal to 2

jimthompson5910 (jim_thompson5910):

f(x) = {x/3, x < or equal to 0 2x - 6, 0 < x < 2 1, x > or equal to 2 literally means f(x) = x/3 if x <= 0 OR f(x) = 2x-6 if 0 < x < 2 OR f(x) = 1 if x >= 2

jimthompson5910 (jim_thompson5910):

To graph this, you graph y = x/3, but you only do so for x values less than or equal to 0. Then you add on y = 2x-6 but only for x values between 0 and 2 (excluding both 0 and 2) Finally, you add on y = 1, and you graph this for values of x greater than or equal to 2

OpenStudy (firejay5):

@jim_thompson5910 How many lines will be graphed

jimthompson5910 (jim_thompson5910):

There are 3 pieces, so 3 lines.

OpenStudy (firejay5):

How would they look on a graph

jimthompson5910 (jim_thompson5910):

just graph y = x/3

jimthompson5910 (jim_thompson5910):

Then erase everything less than x = 0

jimthompson5910 (jim_thompson5910):

do the same with the other pieces

OpenStudy (firejay5):

Are the lines going to be spread out

OpenStudy (firejay5):

@jim_thompson5910 Can you show me all the lines, because I don't get it

jimthompson5910 (jim_thompson5910):

one moment please

OpenStudy (firejay5):

ok

jimthompson5910 (jim_thompson5910):

have a look at this page and tell me if it helps or not http://www.coolmath.com/precalculus-review-calculus-intro/precalculus-algebra/07-graphing-piecewise-defined-functions-01.htm

jimthompson5910 (jim_thompson5910):

it gives a pretty detailed explanation of piecewise functions and how to graph them

OpenStudy (firejay5):

really

OpenStudy (firejay5):

I can graph 2, but don't know how 3 would look

jimthompson5910 (jim_thompson5910):

graph 3 would be a horizontal line at 1 on the y axis

jimthompson5910 (jim_thompson5910):

but you're only graphing from x = 2 and to the right

OpenStudy (firejay5):

graph 3 would be left of y-axis and right of y - axis

jimthompson5910 (jim_thompson5910):

no, just on the right side

jimthompson5910 (jim_thompson5910):

from x = 2 and onward to the right

OpenStudy (firejay5):

@jim_thompson5910 what's the equations for the 1st and 2nd graph and what is the domain and range???

jimthompson5910 (jim_thompson5910):

the equations are the first and second pieces

jimthompson5910 (jim_thompson5910):

then use the graph to find the domain remember the domain is the set of allowed inputs the range is the set of possible outputs

OpenStudy (firejay5):

Range has to be {y/y > 3} for example

jimthompson5910 (jim_thompson5910):

look at the graph and see which y values come out, the complete set or list will make up the range

OpenStudy (firejay5):

Idk

OpenStudy (anonymous):

The domain is all real numbers

OpenStudy (firejay5):

yea I got that

OpenStudy (firejay5):

it's just the rang

OpenStudy (anonymous):

for the range it would probably be easiest to just graph the function and see which y-values are convered

OpenStudy (anonymous):

covered

OpenStudy (firejay5):

What would the y = mx + b look like for all 3

OpenStudy (anonymous):

x/3 = 1/3x+0 2x-6=2x-6 1 = 0x+1 or just a straight, horizontal line from x= 2 to infinity

OpenStudy (firejay5):

so the y values must be the same

OpenStudy (anonymous):

no, because they all have different domains, their ranges will be different

OpenStudy (firejay5):

so how will it work

OpenStudy (firejay5):

what am I looking

OpenStudy (firejay5):

@ArkGoLucky what am I looking for

OpenStudy (anonymous):

sorry was away for a few minutes so the range of a function is all the outputs of the function, or the y values. so graph the functions and see the range of the y values and that's your range

OpenStudy (firejay5):

I got it already sorry man